THE ν_{1} BAND AND MOLECULAR PROPERTIES OF CHLOROSYL FLUORIDE, FClO

\underline{HOLGER S. P. MÜLLER}, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, D-50937 Köln, Germany.

Rotationally resolved spectra of FClO have been obtained for the first time employing the in situ hydrolysis of ClF_{3}. The molecule has also been proposed as an intermediate in several reactions involving ClF, Cl_{2}O, and ClF_{3}O. At present, the ClO stretching band centered near 1037.7 cm^{−1} for F^{35}ClO has been recorded with a resolution of 0.003 cm^{−1}, and its analysis is under way. Strong a-type and much weaker b-type lines have been assigned. Transitions with quantum numbers larger than 70 and 20 for J and K_{a}, respectively, enabled the determination of precise spectroscopic constants. The derived properties, such as molecular structure and force field, will be discussed in relation to ab initio results and data from related molecules, for example FClO_{n}, with n = 0, 2, 3.